# ERDŐS–PÓSA PROPERTY OF CYCLES THAT ARE FAR APART

### v2

@@ 3 The Proof
Let $t:=(e,i,e^{\prime},j)$. Let $P_{0}$, $P_{1}$, and $P_{2}$ be the forest leg, the first leg, and second leg of $W_{t}$, respectively. For the proof of a, recall that $P_{0}$ is a connected subgraph of $F_{0}^{-}$. Since $t$ is admissible, $B_{G}(V(P_{0}),d)$ contains no vertex in $\widehat{Y}$. Since $P_{0}\subseteq F_{0}^{-}$, $V(P_{0})$ contains no vertex of $B_{G}(V(C_{i}),r-d)$ for each $i\in[p]$. Hence, $B_{G}(V(P_{0}),d)$ contains no vertex in $B_{G}(V(C_{i}),r-2d)$ for any $i\in[p]$. Therefore $B_{G}(V(P_{0}),d)=B_{F_{0}}(V(P_{0}),d)$. Since $P_{0}$ is connected in $F_{0}$, $B_{F_{0}}(V(P_{0}),d)$ induces a connected subgraph of $F_{0}$.
Now we prove b. If $\ell\in[p]\setminus\{i,j\}$, then $\Psi_{\ell}(t)$ is empty, so $c_{\ell}=0$. Thus, it is sufficient to prove that $B_{G}(V(P_{1}),d)$ induces a connected subgraph of $F_{i}$ and that $B_{G}(V(P_{2}),d)$ induces a connected subgraph of $F_{j}$. We prove the first of these statements, the proof of the second is symmetric.
Since $P_{1}$ contains no vertex in $Y_{1}\cup\{y_{i}\}$, $P_{1}$ is a connected subgraph of $F_{i}^{-}$. Since $V(F^{-}_{i})\subseteq B_{G}(V(C_{i}),r-d)$, $B_{G}(V(P_{1}),d)\subseteq B_{G}(V(C_{i}),r)$. Furthermore, $B_{G}(V(P_{1}),d)$ contains no vertex in $\widehat{Y}$, so $B_{G}(V(P_{1}),d)=B_{F_{i}}(V(P_{1}),d)$. Since $P_{1}$ is connected in $F_{i}$, $B_{F_{i}}(V(P_{1}),d)$ induces a connected subgraph of $F_{i}$. ∎
@@
We first consider the case in which $v$ is in the forest leg of $W_{t_{1}}$ or $w$ is in the forest leg of $W_{t_{2}}$. Without loss of generality, suppose $v$ is in the forest leg $P_{0}$ of $W_{t_{1}}$. Then $w\in B_{G}(v,d)\subseteq B_{G}(V(P_{0}),d)\subseteq\Psi_{0}(t_{1})$. Now we will show that $w\in\Psi_{0}(t_{2})$. Recall that $w$ lies in one of the three legs of $W_{t_{2}}$. If $w\in V(Q_{0})$ then $w\in V(Q_{0})\subseteq\Psi_{0}(t_{2})$ as desired. Now assume $w$ lies in the first or second leg of $W_{t_{2}}$. Suppose without loss of the generality that $w\in V(Q_{1})$. Let $x$ be the endpoint of the $i$-exit edge $e$ that is contained in $Q_{0}$. Then $w\in B_{G}(x,d)\subseteq\Psi_{0}(t_{2})$. Therefore $\Psi_{0}(t_{1})\cap\Psi_{0}(t_{2})\supseteq\{w\}\supsetneq\emptyset$, a contradiction.
Next we consider the case where $v$ is not in the forest leg of $W_{t_{1}}$ and $w$ is not in the forest leg of $W_{t_{2}}$. Without loss of generality, suppose $v$ is in the first leg $P_{1}$ of $W_{t_{1}}$ and $w$ is in the first leg $Q_{1}$ of $W_{t_{2}}$. Recall that $t_{2}=(e,\ell,e^{\prime},\ell^{\prime})$, so $Q_{1}\subseteq F_{\ell}^{-}$. Then $v\in B_{G}(w,d)\subseteq B_{G}(V(Q_{1}),d)\subseteq\Psi_{\ell}(t_{2})$. Since $t_{2}$ is admissible and $w\in V(F^{-}_{\ell})$ we have that $B_{G}(w,d)\subseteq V(F_{\ell})$ and so $v\in V(F_{\ell})$. Since $v\not\in V(F_{0}^{-})$ and $v\not\in\widehat{Y}$, $v\in B_{G}(V(C_{j}),r-d)$ for some $j\in\{1,\ldots,p\}$. Since $v\notin B_{G}(V(C_{\ell}),r)\cap B_{G}(V(C_{j}),r)$ for any $j\neq\ell$, it must be that $j=\ell$. Thus $v\in V(F^{-}_{\ell})$. Hence $v\in\Psi_{\ell}(t_{1})$. Therefore, $\Psi_{\ell}(t_{1})\cap\Psi_{\ell}(t_{2})\supseteq\{v\}\supsetneq\emptyset$, a contradiction. ∎
@@
Note that $\Psi_{\ell}(t)$ is non-empty only for the values of $\ell\in\{0,i,j\}$. If $\ell=0$ then, by definition, $X$ contains a vertex in $B_{G}(V(P_{0}),d)=\Psi_{0}(t)$. Therefore $B_{G}(X,d)\subseteq B_{G}(X,r)$ contains a vertex of $P_{0}$ and there is nothing more to prove. Otherwise, let $P_{1}$ be the first leg of $t$ and let $P_{2}$ be the second leg of $t$. If $\ell=i$ then $\operatorname{dist}_{G}(X,V(P_{1}))\leq d$. Therefore $B_{G}(X,d)$ contains a vertex of $P_{1}$. Therefore $B_{G}(X,r)$ contains all vertices of $P_{1}$, including the endpoint of $e$ in $P_{1}$. If $\ell=j$, then the same argument shows that $B_{G}(X,r)$ contains the endpoint of $e^{\prime}$ in $P_{2}$. ∎
Let $\leavevmode\color[rgb]{0.76,0.13,0.28}F^{\star}$ be the forest obtained by taking disjoint copies $\leavevmode\color[rgb]{0.76,0.13,0.28}F_{0}^{\star},\ldots,F_{p}^{\star}$ of $F_{0},\ldots,F_{p}$, respectively. That is, $F_{0}^{\star},\ldots,F_{p}^{\star}$ are pairwise vertex-disjoint and $F_{i}^{\star}$ is isomorphic to $F_{i}$, for each $i\in\{0,\ldots,p\}$. For each $i\in\{0,\ldots,p\}$ and each $v\in V(F_{i})$, define $\leavevmode\color[rgb]{0.76,0.13,0.28}v_{i}$ to be the unique vertex of $F^{\star}_{i}$ that corresponds to $v$.
Let $t$ be an admissible tuple. Let $\leavevmode\color[rgb]{0.76,0.13,0.28}\Psi^{\star}(t):=\bigcup_{i=0}^{p}\{v_{i}:v\in\Psi_{i}(t)\}$. By Claim 10, $F[\Psi^{\star}(t)]$ has at most three components. Let
$\leavevmode\color[rgb]{0.76,0.13,0.28}\mathcal{A}:=\{\Psi^{\star}(t):\text{$t$is an admissible tuple}\}.$
By Theorem 3 applied for $\leavevmode\color[rgb]{0.76,0.13,0.28}k^{\star}:=k+\tfrac{1}{2}s(k)$, $c=3$, $F$, and $\mathcal{A}$, at least one of the following is true: (a) there are $k^{\star}$ pairwise vertex-disjoint members of $\mathcal{A}$; or (b) there exists $X^{\star}\subseteq V(F)$ with $|X^{\star}|\leq\ell^{\star}(k^{\star},3)$ and such that $X^{\star}\cap A\neq\emptyset$ for each $A\in\mathcal{A}$.
First, we consider outcome Item b and let $X^{\star}$ be the promised subset of $V(F^{\star})$. Define
$\leavevmode\color[rgb]{0.76,0.13,0.28}X_{3}:=\{v\in V(G):v_{i}\in X^{\star}\textrm{ for some $i\in\{0,\ldots,p\}$}\}.$
Note that for every admissible tuple $t$,
@@
$\leavevmode\color[rgb]{0.76,0.13,0.28}X:=\widehat{X}\cup X_{3}\enspace.$
@@
$|X|=|\widehat{X}|+|X_{3}|<\textstyle 2k+2k^{2}+(\binom{k}{2}+k)\cdot s(k)+\ell^{\star}(k+\tfrac{1}{2}s(k),3)=f(k).$
@@
Consider an arbitrary cycle $C$ in $G$. By Claim 9, either $C$ contains a vertex in $B_{G}(\widehat{Y},r)$ or $C$ contains the extended forest leg of some admissible tuple $t$. In the former case, by (6), $\widehat{Y}\subseteq B_{G}(\widehat{X},13d)$ so $B_{G}(\widehat{Y},r)\subseteq B_{G}(\widehat{X},19d)$ and we are done. Therefore, we can assume that $C$ contains the extended forest leg of $W_{t}$ for some admissible tuple $t$. By (8) and Claim 12, $B_{G}(X_{3},r)$ contains a vertex of $C$.
Now we consider outcome Item a: there are $k^{\star}$ pairwise vertex-disjoint members $\leavevmode\color[rgb]{0.76,0.13,0.28}A_{1},\ldots,A_{k^{\prime}}$ of $\mathcal{A}$. For each $j\in[k^{\star}]$, let $\leavevmode\color[rgb]{0.76,0.13,0.28}t_{j}$ be an admissible tuple such that $A_{j}=\Psi^{\star}(t_{j})$. Let $\leavevmode\color[rgb]{0.76,0.13,0.28}T:=\{t_{1},\ldots,t_{k^{\prime}}\}$. Therefore, for every $i\in\{0,\ldots,p\}$ and every $t,t^{\prime}\in T$ with $t\neq t^{\prime}$ we have $\Psi_{i}(t)\cap\Psi_{i}(t^{\prime})=\emptyset$. By Claim 11,
@@
Recall that, for each $t\in T$, $W_{t}$ is a path in $G$ or contains a cycle in $G$. Let $\leavevmode\color[rgb]{0.76,0.13,0.28}J:=\{t\in T:\text{$W_{t}$ contains a cycle}\}$. By (9) the cycles defined by $W_{t}$ for $t\in J$ define a $d$-packing of cycles in $G$. Therefore $|J|<k$ and $|T\setminus J|>k^{\star}-k=\tfrac{1}{2}s(k)$.
Let $\mathcal{C}^{\prime}$ be the subset of $\{C_{1},\ldots,C_{p}\}$ that contains each $C_{i}$ if and only $V(C_{i})$ contains an endpoint of $W_{t}$ for some $t\in T\setminus J$. Define the graph $G^{\prime}:=\bigcup_{C\in\mathcal{C}^{\prime}}C\cup\bigcup_{t\in T\setminus J}W_{t}$. Note that all vertices of $G^{\prime}$ are of degree $2$ or $3$. Since $G^{\prime}$ does not contain $W_{t}$ for any $t\in J$, each degree $3$ vertex of $G^{\prime}$ is an endpoint of $W_{t}$ for some $t\in T\setminus J$. Therefore, $G^{\prime}$ contains $2|T\setminus J|\geq s(k)$ vertices of degree $3$. By Theorem 2, $G^{\prime}$ contains a set $\mathcal{D}$ of $k$ pairwise vertex-disjoint cycles. The degree-$3$ vertices of $G^{\prime}$ partition the edges of each cycle in $\mathcal{D}$ into paths, where each such path is either contained in a cycle $C_{i}$ for some $i\in[p]$ or is a path $W_{t}$ for some $t\in T\setminus J$.
We now show that $\mathcal{D}$ is a $d$-packing of cycles in $G$. Let $D$ and $D^{\prime}$ be two distinct cycles in $\mathcal{D}$. Let $P$ be a shortest path in $G$ from $V(D)$ to $V(D^{\prime})$. Let $v\in V(D)$ and $v^{\prime}\in V(D^{\prime})$ be the endpoints of $P$. If $v\in W_{t}$ and $v^{\prime}\in W_{t^{\prime}}$ for some $t,t^{\prime}\in T$ such that $W_{t}\subseteq D$ and $W_{t^{\prime}}\subseteq D^{\prime}$ then, since $D$ and $D^{\prime}$ are vertex-disjoint we have $t\neq t^{\prime}$, and by (9) we have $\operatorname{dist}_{G}(v,v^{\prime})\geq\operatorname{dist}_{G}(V(W_{t}),V(W_{t}))>d$, as desired.
@@
Since $v^{\prime}\in B_{G}(V(C_{i}),d)\subseteq B_{G}(V(C_{i}),r-d)$, $v^{\prime}$ is not in the forest leg of $W_{t^{\prime}}$. Assume, without loss of generality, that $v^{\prime}$ is in the first leg $P_{0}^{\prime}$ of $W_{t^{\prime}}$. Since $t^{\prime}$ is admissible, $P_{0}^{\prime}$ contains no vertex of $Y_{1}\supseteq\bigcup_{j\in[p]\setminus\{i\}}(B_{G}(V(C_{i}),r)\cap B_{G}(V(C_{j}),r)$. Therefore, $P_{0}^{\prime}$ connects a vertex of $V(C_{i})$ with an $i$-exit edge. Since $C_{i}$ is $d$-unicyclic and the first edge of $P$ is not an edge of $C_{i}$, $P_{0}^{\prime}$ contains the first edge of $P$. Therefore, $P\subseteq P_{0}^{\prime}\subseteq W_{t^{\prime}}\subseteq D^{\prime}$. Since $v\in V(D)$ is incident to the first edge of $P$, this contradicts the fact that $D$ and $D^{\prime}$ are vertex disjoint. This completes the proof that $\mathcal{D}$ is a $d$-packing in $G$. Since $|\mathcal{D}|\geq k$ this is the final contradiction that completes the proof of Theorem 8. ∎
4 Discussion
We conjecture that Theorem 1 remains true with $f(k)\in\mathcal{O}(k\log k)$ as in the original Erdős-Pósa theorem, which would be optimal. Our proof gives $f(k)\in\mathcal{O}(k^{18}\operatorname{polylog}k)$. By applying the techniques in the appendix directly to the tuples $\{(\Psi_{\ell}(t))_{\ell\in\{0,\ldots,p\}}:\text{$t$ is an admissible tuple}\}$ that appear in the proof of Theorem 8, the exponent $18$ could be reduced with some additional work but would still be much greater than $1$.
Turning to the function $g(d)$ in Theorem 1, we remark that we made no attempt to minimize the leading constant in our proof, which shows $g(d)=19d$. We believe that we can decrease the constant $19$ somewhat, at the price of a more complicated proof, but it would still be above $10$. It would be interesting to determine the smallest function $g(d)$ for which Theorem 1 is true (with no constraint on $f(k)$).
Acknowledgements
The authors would like to thank Ugo Giocanti, Jędrzej Hodor, Freddie Illingworth, and Clément Legrand-Duchesne for discussions in the early stage of this research. This research was initiated at the Third Ottawa-Carleton Workshop on Graphs and Probability, held in Mont Sainte-Marie, Québec and in Toronto, Ontario, Oct 13–20, 2024.
Appendix

### v3

@@ 4 The Proof
Let $t:=(e,i,e^{\prime},j)$. Let $P_{0}$, $P_{1}$, and $P_{2}$ be the forest leg, the first leg, and second leg of $W_{t}$, respectively. For the proof of ˜a, recall that $P_{0}$ is a connected subgraph of $F_{0}^{-}$. Since $t$ is admissible, $B_{G}(V(P_{0}),d)$ contains no vertex in $\widehat{Y}$. Since $P_{0}\subseteq F_{0}^{-}$, $V(P_{0})$ contains no vertex of $B_{G}(V(C_{i}),r-d)$ for each $i\in[p]$. Hence, $B_{G}(V(P_{0}),d)$ contains no vertex in $B_{G}(V(C_{i}),r-2d)$ for any $i\in[p]$. Therefore $B_{G}(V(P_{0}),d)=B_{F_{0}}(V(P_{0}),d)$. Since $P_{0}$ is connected in $F_{0}$, $B_{F_{0}}(V(P_{0}),d)$ induces a connected subgraph of $F_{0}$.
Now we prove ˜b. If $\ell\in[p]\setminus\{i,j\}$, then $\Psi_{\ell}(t)$ is empty, so $c_{\ell}=0$. Assume thus that $\ell\in\{i,j\}$. We prove that $B_{G}(V(P_{1}),d)$ induces a connected subgraph of $F_{i}$ and that $B_{G}(V(P_{2}),d)$ induces a connected subgraph of $F_{j}$. By the definition of $\Psi_{\ell}(t)$, this implies that $c_{\ell}=1$ in case $i\neq j$, and $c_{\ell}\leq 2$ in case $i=j$. We prove the first of these statements, the proof of the second is symmetric.
Since $P_{1}$ contains no vertex in $Y_{1}\cup\{y_{i}\}$, $P_{1}$ is a connected subgraph of $F_{i}^{-}$. Since $V(F^{-}_{i})\subseteq B_{G}(V(C_{i}),r-d)$, $B_{G}(V(P_{1}),d)\subseteq B_{G}(V(C_{i}),r)$. Furthermore, $B_{G}(V(P_{1}),d)$ contains no vertex in $\widehat{Y}$ as $t$ is admissible, so $B_{G}(V(P_{1}),d)=B_{F_{i}}(V(P_{1}),d)$. Since $P_{1}$ is connected in $F_{i}$, $B_{F_{i}}(V(P_{1}),d)$ induces a connected subgraph of $F_{i}$. ∎
@@
We first consider the case in which $v$ is in the forest leg of $W_{t_{1}}$ or $w$ is in the forest leg of $W_{t_{2}}$. Without loss of generality, suppose $v$ is in the forest leg $P_{0}$ of $W_{t_{1}}$. Then $w\in B_{G}(v,d)\subseteq B_{G}(V(P_{0}),d)\subseteq\Psi_{0}(t_{1})$. Now we will show that $w\in\Psi_{0}(t_{2})$. Recall that $w$ lies in one of the three legs of $W_{t_{2}}$. If $w\in V(Q_{0})$ then $w\in V(Q_{0})\subseteq\Psi_{0}(t_{2})$ as desired. Now assume $w$ lies in the first or second leg of $W_{t_{2}}$. Suppose without loss of the generality that $w\in V(Q_{1})$. Let $x$ be the endpoint of the $\ell$-exit edge $e$ that lies in $Q_{0}$. Observe that the ball $B_{G}(w,d)$ hits $F^{-}_{0}$, since it contains $v$. By the definition of $Q_{1}$, this implies that $x\in B_{G}(w,d)$. But then $w\in B_{G}(x,d)\subseteq\Psi_{0}(t_{2})$, and therefore $\supsetneq\Psi_{0}(t_{1})\cap\Psi_{0}(t_{2})\supseteq\{w\}\supsetneq\emptyset$, a contradiction.
Next we consider the case where $v$ is not in the forest leg of $W_{t_{1}}$ and $w$ is not in the forest leg of $W_{t_{2}}$. Without loss of generality, suppose $v$ is in the first leg $P_{1}$ of $W_{t_{1}}$ and $w$ is in the first leg $Q_{1}$ of $W_{t_{2}}$. Recall that $t_{2}=(e,\ell,e^{\prime},\ell^{\prime})$, so $Q_{1}\subseteq F_{\ell}^{-}$. Since $\operatorname{dist}_{G}(v,w)\leq d$ by assumption, we then have $v\in B_{G}(w,d)\subseteq B_{G}(V(Q_{1}),d)\subseteq\Psi_{\ell}(t_{2})$. Since $t_{2}$ is admissible and $w\in V(F^{-}_{\ell})$ we have that $B_{G}(w,d)\subseteq V(F_{\ell})$ and so $v\in V(F_{\ell})$. Since $v\not\in V(F_{0}^{-})$ and $v\not\in\widehat{Y}$, $v\in B_{G}(V(C_{j}),r-d)$ for some $j\in\{1,\ldots,p\}$. Since $v\notin B_{G}(V(C_{\ell}),r)\cap B_{G}(V(C_{j}),r)$ for any $j\neq\ell$, it must be that $j=\ell$. Thus $v\in V(F^{-}_{\ell})$. Hence $v\in\Psi_{\ell}(t_{1})$. Therefore, $\supsetneq\Psi_{\ell}(t_{1})\cap\Psi_{\ell}(t_{2})\supseteq\{v\}\supsetneq\emptyset$, a contradiction. ∎
@@
Note that $\Psi_{\ell}(t)$ is non-empty only for the values of $\ell\in\{0,i,j\}$. If $\ell=0$ then, by assumption, $X$ contains a vertex in $B_{G}(V(P_{0}),d)=\Psi_{0}(t)$. Therefore $B_{G}(X,d)\subseteq B_{G}(X,r)$ contains a vertex of $P_{0}$ and there is nothing more to prove. Otherwise, let $P_{1}$ be the first leg of $W_{t}$ and let $P_{2}$ be the second leg of $W_{t}$. If $\ell=i$ then $\operatorname{dist}_{G}(X,V(P_{1}))\leq d$. Therefore $B_{G}(X,d)$ contains a vertex of $P_{1}$. Therefore $B_{G}(X,r)$ contains all vertices of $P_{1}$, including the endpoint of $e$ in $P_{1}$. If $\ell=j$, then the same argument shows that $B_{G}(X,r)$ contains the endpoint of $e^{\prime}$ in $P_{2}$. ∎
Let $\color[rgb]{0.76,0.13,0.28}F^{\star}$ be the forest obtained by taking disjoint copies $\color[rgb]{0.76,0.13,0.28}F_{0}^{\star},\ldots,F_{p}^{\star}$ of $F_{0},\ldots,F_{p}$, respectively. That is, $F_{0}^{\star},\ldots,F_{p}^{\star}$ are pairwise vertex-disjoint and $F_{i}^{\star}$ is isomorphic to $F_{i}$, for each $i\in\{0,\ldots,p\}$. For each $i\in\{0,\ldots,p\}$ and each $v\in V(F_{i})$, define $\color[rgb]{0.76,0.13,0.28}v_{i}$ to be the unique vertex of $F^{\star}_{i}$ that corresponds to $v$.
Let $t$ be an admissible tuple. Let $\color[rgb]{0.76,0.13,0.28}\Psi^{\star}(t):=\bigcup_{i=0}^{p}\{v_{i}:v\in\Psi_{i}(t)\}$. By Claim˜10, $F^{\star}[\Psi^{\star}(t)]$ has at most three components. Let
$\color[rgb]{0.76,0.13,0.28}\mathcal{A}:=\{F^{\star}[\Psi^{\star}(t)]:\text{$t$ is an admissible tuple}\}.$
By Theorem˜3 applied for $\color[rgb]{0.76,0.13,0.28}k^{\star}:=k+\tfrac{1}{2}s(k)$, $c=3$, $F^{\star}$, and $\mathcal{A}$, at least one of the following is true: (a) there are $k^{\star}$ pairwise vertex-disjoint members of $\mathcal{A}$; or (b) there exists $X^{\star}\subseteq V(F^{\star})$ with $|X^{\star}|\leq\ell^{\star}(k^{\star},3)$ and such that $X^{\star}\cap V(A)\neq\emptyset$ for each $A\in\mathcal{A}$.
First, we consider outcome Item˜b and let $X^{\star}$ be the promised subset of $V(F^{\star})$. Define
$\color[rgb]{0.76,0.13,0.28}X_{3}:=\{v\in V(G):v_{i}\in X^{\star}\textrm{ for some $i\in\{0,\ldots,p\}$}\}.$
Note that for every admissible tuple $t$, $X^{\star}$ intersects $\Psi^{\star}(t)$, and therefore
@@
$\color[rgb]{0.76,0.13,0.28}X:=\widehat{X}\cup X_{3}.$
@@
$|X|\leq|\widehat{X}|+|X_{3}|<\textstyle 2k+2k^{2}+(\binom{k}{2}+k)\cdot s(k)+\ell^{\star}(k+\tfrac{1}{2}s(k),3)=f(k).$
@@
Consider an arbitrary cycle $C$ in $G$. By Claim˜9, either $C$ contains a vertex in $B_{G}(\widehat{Y},r)$ or $C$ contains the extended forest leg of some admissible tuple $t$. In the former case, by (6), $\widehat{Y}\subseteq B_{G}(\widehat{X},13d)$ so $B_{G}(\widehat{Y},r)\subseteq B_{G}(\widehat{X},19d)$ and we are done. Therefore, we can assume that $C$ contains the extended forest leg of $W_{t}$ for some admissible tuple $t$. By (8) and Claim˜12, $B_{G}(X_{3},r)$ contains a vertex of $C$.
Now we consider outcome Item˜a: there are $k^{\star}$ pairwise vertex-disjoint members $\color[rgb]{0.76,0.13,0.28}A_{1},\ldots,A_{k^{\star}}$ of $\mathcal{A}$. For each $j\in[k^{\star}]$, let $\color[rgb]{0.76,0.13,0.28}t_{j}$ be an admissible tuple such that $A_{j}=F^{\star}[\Psi^{\star}(t_{j})]$. Let $\color[rgb]{0.76,0.13,0.28}T:=\{t_{1},\ldots,t_{k^{\star}}\}$. Therefore, for every $i\in\{0,\ldots,p\}$ and every $t,t^{\prime}\in T$ with $t\neq t^{\prime}$ we have $\Psi_{i}(t)\cap\Psi_{i}(t^{\prime})=\emptyset$. By Claim˜11,
@@
Recall that, for each $t\in T$, $W_{t}$ is a path in $G$ or contains a cycle in $G$. Let $\color[rgb]{0.76,0.13,0.28}J:=\{t\in T:\text{$W_{t}$ contains a cycle}\}$. By (9) the cycles defined by $W_{t}$ for $t\in J$ define a $d$-packing of cycles in $G$. Therefore $|J|<k$ and $|T\setminus J|>k^{\star}-k=\tfrac{1}{2}s(k)$.
Let $\mathcal{C}^{\prime}$ be the subset of $\{C_{1},\ldots,C_{p}\}$ that contains each $C_{i}$ if and only if $V(C_{i})$ contains an endpoint of $W_{t}$ for some $t\in T\setminus J$. Define the graph $G^{\prime}:=\bigcup_{C\in\mathcal{C}^{\prime}}C\cup\bigcup_{t\in T\setminus J}W_{t}$. Recall that the cycles of $\{C_{1},\ldots,C_{p}\}$ are vertex disjoint by choice, and that all $W_{t}$’s with $t\in T$ are vertex disjoint by 9. Thus all vertices of $G^{\prime}$ are of degree $2$ or $3$. Since $G^{\prime}$ does not contain $W_{t}$ for any $t\in J$, each degree $3$ vertex of $G^{\prime}$ is an endpoint of $W_{t}$ for some $t\in T\setminus J$. Therefore, $G^{\prime}$ contains $2|T\setminus J|\geq s(k)$ vertices of degree $3$. By Theorem˜2, $G^{\prime}$ contains a set $\mathcal{D}$ of $k$ pairwise vertex-disjoint cycles. The degree-$3$ vertices of $G^{\prime}$ partition the edges of each cycle in $\mathcal{D}$ into paths, where each such path is either contained in a cycle $C_{i}$ for some $i\in[p]$ or is a path $W_{t}$ for some $t\in T\setminus J$.
We now show that $\mathcal{D}$ is a $d$-packing of cycles in $G$. Let $D$ and $D^{\prime}$ be two distinct cycles in $\mathcal{D}$. Let $P$ be a shortest path in $G$ from $V(D)$ to $V(D^{\prime})$. Let $v\in V(D)$ and $v^{\prime}\in V(D^{\prime})$ be the endpoints of $P$. If $v\in W_{t}$ and $v^{\prime}\in W_{t^{\prime}}$ for some $t,t^{\prime}\in T$ such that $W_{t}\subseteq D$ and $W_{t^{\prime}}\subseteq D^{\prime}$ then, since $D$ and $D^{\prime}$ are vertex-disjoint we have $t\neq t^{\prime}$, and by (9) we have $\operatorname{dist}_{G}(v,v^{\prime})\geq\operatorname{dist}_{G}(V(W_{t}),V(W_{t^{\prime}}))>d$, as desired.
@@
Since $^{\prime}\in B_{G}(V(C_{i}),d)\subseteq B_{G}(V(C_{i}),r-d), vv^{\prime}$ is not in the forest leg of $W_{t^{\prime}}$. Assume, without loss of generality, that $v^{\prime}$ is in the first leg $P_{1}^{\prime}$ of $W_{t^{\prime}}$. Since $t^{\prime}$ is admissible and by (3), $P_{1}^{\prime}$ contains no vertex of $\bigcup_{j\in[p]\setminus\{i\}}(B_{G}(V(C_{i}),r)\cap B_{G}(V(C_{j}),r)$. Therefore, $P_{1}^{\prime}$ connects a vertex of $V(C_{i})$ with an $i$-exit edge. Since $C_{i}$ is $d$-unicyclic and the first edge of $P$ is not an edge of $C_{i}$, $P_{1}^{\prime}$ contains the first edge of $P$. Therefore, $P\subseteq P_{1}^{\prime}\subseteq W_{t^{\prime}}\subseteq D^{\prime}$. Since $v\in V(D)$ is incident to the first edge of $P$, this contradicts the fact that $D$ and $D^{\prime}$ are vertex disjoint. This completes the proof that $\mathcal{D}$ is a $d$-packing in $G$. Since $|\mathcal{D}|\geq k$ this is the final contradiction that completes the proof of Theorem˜8. ∎
5 Proof of Theorem˜3